# RATE: Score Reward Models with Imperfect Rewrites of Rewrites

### v1

@@ 4 Theoretical Analysis of RATE
    This means that we don’t need to worry about potential interactions between rewrite errors (affecting $\xi$) and other attributes of the response ($Z$), even if $W$ and $Z$ have interactions.
@@
    We assume that the off-target changes introduced by the rewrite process are randomly drawn from a distribution determined by the particular rewrite method being used. That is,
[…]
    For example, when our rewriter is GPT-4o, the off-target yet mutable attributes such as specific word choice and grammatical structure are drawn from a ‘GPT-like’ distribution.
Theorem 1 (Unbiasedness and Consistency).
[…]
$\hat{\tau}_{\text{ATT}}$ $=\frac{1}{n_{1}}\sum_{(i,j):w^{ij}=1}[R(x^{i},\text{Re}(\text{Re}(x^{i},y^{ij},0),1))-R(x^{i},\text{Re}(x^{i},y^{ij},0))]$
$\hat{\tau}_{\text{ATU}}$ $=\frac{1}{n_{0}}\sum_{(i,j):w^{ij}=0}[R(x^{i},\text{Re}(x^{i},y^{ij},1))-R(x^{i},\text{Re}(\text{Re}(x^{i},y^{ij},1),0))]$
$\hat{\tau}_{\text{ATE}}$ $=\frac{n_{1}}{n_{0}+n_{1}}\hat{\tau}_{\text{ATT}}+\frac{n_{0}}{n_{0}+n_{1}}\hat{\tau}_{\text{ATU}}$
where $n_{1}$ and $n_{0}$ are the number of pairs with observed $W=1$ and $W=0$ respectively, are unbiased and consistent estimators of the ATT, ATU, and ATE.
See Appendix A for the proof.
5 Experiments
We evaluate reward models using RATE on real-world and synthetic data. Experiments show:
  • Across a variety of attributes and datasets, RATE gives substantively different estimates compared to the naive (non-causal) baseline.
  • In semi-synthetic data with known ground truth behavior, RATE is robust to distributional shift, while the naive estimator is not.
  • Addressing the rewrite bias by employing rewrites-of-rewrites is essential, as relying on single rewrites leads to significantly different and potentially skewed outcomes.
Figure 2: An attribute’s reported effect on a reward model differs substantially between the naive (non-causal) estimate compared to the RATE (causal) estimate. The naive estimator overstates the length bias of FsfairX-LLaMA3-RM-v0.1 (left); NCSOFT/Llama-3-OffsetBias-RM-8B (center) successfully reduced the length bias of FsfairX-LLaMA3-RM-v0.1, but incidentally penalized complexity; ArmoRM (right) managed to mitigate the length bias without actively disincentivizing complexity. Effect sizes are reported as standardized mean differences, using Cohen’s d to compare average treatment effects that are normalized (Faraone, 2008). Bars represent a 95% confidence interval.
Real World Reward Models
We select several of the top-performing reward models from RewardBench (Lambert et al., 2024) and evaluate them using both RATE and the naive method across a variety of attributes and datasets: IMDB (Maas et al., 2011), ELI5 (Fan et al., 2019), HelpSteer (Wang et al., 2023). Randomly sampled rewrites with associated rewards are shown in Appendix B, along with details for designing rewrite instructions.
Figure 2 shows the estimated response of each reward model to each attribute. Of particular interest are the evaluations of FsfairX-LLaMA3-RM-v0.1 (Dong et al., 2023) and NCSOFT (Park et al., 2024a) with respect to length. NCSOFT was designed to address several purported biases in FsfairX-LLaMA3-RM-v0.1, including length. The contrast between RATE and the naive estimate suggests that the length bias for FsfairX-LLaMA3-RM-v0.1 may have been overreported due to non-causal correlations in evaluation. At any rate, NCSOFT successfully removed the remaining length bias.
Synthetic Experiments
While the real-world experiments demonstrate RATE’s practical utility, they don’t allow us to verify its accuracy against a known ground truth. To address this, we turn to synthetic experiments where we can control the underlying data generation process and introduce known correlations between attributes. See Appendix B for details.
Is RATE correctly capturing the ATE? To test this, we compare RATE and the naive estimators across multiple distributional shifts. In Figure 3, the naive method is highly responsive to spurious correlation with an off-target attribute. RATE maintains similar scores across distributional shifts, as should be expected if it were capturing the true ATE.
In Figure 3 (left) we use a DistilBERT sentiment classifier (Socher et al., 2013; Sanh et al., 2020) as a reward model with a ground-truth ATE assumed to be near-zero. Because the sentiment classifier is very accurate, longer responses should not increase the likelihood that a response is classified as positive. We then introduce a correlation between response length and positive sentiment (see Table 6), and show that the naive estimator shows a large effect size. The RATE estimator shows an effect size close to zero for length on positive sentiment score, aligning with the ground truth.
In fig. 3 (right), we evaluate ArmoRM (Wang et al., 2024a) in a similar manner on the HelpSteer dataset. Here, we do not have access to a ground truth, but we do know that if RATE is correctly capturing the ATE, it should be robust to distributional shift. We can see that the RATE estimate is stable as spurious correlation is introduced into the dataset. The naive estimator, on the other hand, is highly sensitive to this correlation, suggesting that it is not capturing the true ATE.
Figure 3: The RATE estimator is robust to distributional shift and better approximates the (assumed) near-zero ATE of length on DistilBERT. Sample size = 9374 for all levels of correlation for the IMDB experiment, and 5148 for the HelpSteer experiment. 95% confidence intervals are shown.
Rewrites of Rewrites vs. Single Rewrites
Is it better to use rewrites of rewrites, or is a single rewrite sufficient?
RATE uses rewrites of rewrites to estimate the causal effect of an attribute on a reward model, addressing potential biases introduced by the rewrite process. Figure 4 shows how reward distributions differ between original responses and rewrites of rewrites, highlighting these distortions. Note that these distortions are not always favorable; while rewrites often correct formatting and make text more ‘GPT-like,’ increasing rewards as in Table 3, they can also produce odd completions. For instance, GPT-4o changed “always the same size" to “annoyingly the same size" when rewriting negative sentiment (see Table 5).
Prompt Original (W = 0) Rewrite of Rewrite (W = 0)
How significant are these distortions? Figure 5 illustrates that the ‘double rewrite’ method produces substantially different estimates compared to the ‘single rewrite’ method. In this case, we intervene on the length attribute in the ELI5 dataset, corresponding to the distortions shown in Figure 4 (right). Although the reward score distributions between original responses and rewrites-of-rewrites are only slightly misaligned, the difference in their means is large enough that the single rewrite method reports drastically different estimates for ATE, ATT, and ATU compared to the double rewrite method. This is not unique to the (Length, ELI5) pair; we observe similar discrepancies across multiple attributes and datasets (see Appendix B).
Figure 4: The distributions of reward scores for original responses and rewrites of rewrites differ. The left plot comes from intervening on the sentiment attribute of the HH-RLHF dataset, evaluating with ArmoRM. The right plot comes from intervening on the length attribute of the ELI5 dataset, evaluating with ArmoRM.
Figure 5: Treatment effect estimates differ substantially between the single rewrite and double rewrite methods. Bars represent a 95% confidence interval.
Implementation Details
@@ 5 Experiments
For all experiments, we use OpenAI BatchAPI to generate rewrites of text, instructing the LLM to modify the target attribute without changing any other aspects of the response (see Table 2). We use the ‘gpt-4o-2024-08-06’ model, incurring \$1.25 per 1M input tokens and \$5.00 per 1M output tokens. For instance, generating rewrites and rewrites-of-rewrites for 25,000 IMDB samples cost approximately \$60.
An important limitation of our implementation is that our chosen rewrite method does not actually use the prompt in the rewrite process. Though this may not be a problem for attributes like sentiment or length, it could be an issue for more complex attributes like helpfulness. We chose this method for its simplicity and ease of use, though future work could explore more sophisticated methods that incorporate the prompt.
Whether or not the prompt is included, crafting instructions to generate appropriate rewrites requires examining rewritten examples and adjusting the instructions accordingly to account for unexpected behavior. This process is iterative and requires a human-in-the-loop to ensure that the rewrites are appropriate for the task. In particular, safety-tuned LLMs are reluctant to rewrite text to be more unhelpful, and so the completions must be carefully examined to ensure that the LLM is willing to generate the desired rewrites.
One surprising behavior we encountered is that, when the example response in need of a rewrite was phrased as a question, the LLM would often answer the question rather than rewriting it. Based on this, we included explicit instructions not to answer questions but, rather, to rewrite them for the HH-RLHF dataset.
6 Discussion
Generalization to Contrastive Rewards
The RATE procedure applies more generally to contrastive rewards of the form $R(x,y_{1},y_{0})$, which assign a relative reward for $y_{1}$ compared to $y_{0}$. In this case, RATE enables us to compute $\mathbb{E}[R(X,Y(W=1),Y(W=0))]$, the expected increase in relative reward attributable to changing attribute $W$ in isolation of everything else, simply by replacing the summands in the earlier formulations. Specifically, we can modify our RATE estimators as follows:
$\hat{\tau}_{\text{ATT}}$ $=\frac{1}{n_{1}}\sum_{(i,j):w^{ij}=1}[R(x^{i},\text{Re}(\text{Re}(x^{i},y^{ij},0),1),\text{Re}(x^{i},y^{ij},0))]$
$\hat{\tau}_{\text{ATU}}$ $=\frac{1}{n_{0}}\sum_{(i,j):w^{ij}=0}[R(x^{i},\text{Re}(x^{i},y^{ij},1),\text{Re}(\text{Re}(x^{i},y^{ij},1),0))]$
$\hat{\tau}_{\text{ATE}}$ $=\frac{n_{1}}{n_{0}+n_{1}}\hat{\tau}_{\text{ATT}}+\frac{n_{0}}{n_{0}+n_{1}}\hat{\tau}_{\text{ATU}}$
As an example of why this may be useful, consider an evaluator LLM that takes a prompt and two responses and returns a preference for which is better. We may view this as a contrastive reward with outputs in ${0,1}$. RATE enables us to estimate how sensitive this evaluator is to different attributes considered in isolation. Notice that in the particular case where $R(x,y_{1},y_{0})=R(x,y_{1})-R(x,y_{0})$, as in, e.g., the Bradley-Terry model, the contrastive RATE estimate is the same as the pointwise RATE estimate described in the main body of the paper. This highlights the versatility of our approach, as it naturally extends to both pointwise and contrastive reward models.
Dynamic Benchmarking
Static benchmarks offer limited insight for model deployment compared to dynamic benchmarking, which is less vulnerable to memorization and can be easily tailored to specific task constraints (Saxon et al., 2024). While the evaluations in this work augment static datasets for the sake of demonstrating its validity, RATE can be easily adapted to dynamic benchmarking by rewriting responses in real-time.
Rewriting the Prompt
Wang et al. (2024b) showed that rewriting prompts outperforms rewriting completions when generating synthetic preference data. Though applied to generic preferences (rather than specific attributes), this suggests that rewriting the prompt may be a useful extension of our method. That is, we could rewrite the prompt to change the attribute of interest, and then generate a completion as usual (the same for rewrites of rewrites). Further research in this direction would need to adapt the latent variable model and consequent RATE estimator, but it could be a promising direction for future work.
Beyond Binary Concepts
This paper focuses on binary attributes, in line with binary treatments in causal inference. Although this may seem limiting, continuous attributes like length can be binarized using thresholds (e.g., above or below a character count), and categorical attributes can be simplified with binary contrasts. This approach works well for many applications, but future work could explore explicit handling of continuous and categorical attributes.
@@ 7 Related Work
Our work intersects with three main areas of research: challenges in reward modeling, causal inference applied to text classifiers, and the use of counterfactuals in language models.
Challenges in Reward Modeling
Our work is particularly motivated by the challenges identified in reward modeling. Lambert et al. (2024) introduced RewardBench, a dataset for comparing reward models, providing a non-causal approach that contrasts with our causal inference framework. Casper et al. (2023) highlighted issues such as misgeneralization and reward hacking in reward models, which our work addresses by quantifying how reward models incentivize specific attributes. Gleave et al. (2021) offered a global metric for comparing reward models, while our approach provides a more fine-grained analysis focused on specific attributes.
Causal Inference Applied to Text Classifiers
Much previous work on causal inference in NLP focuses on text classifiers, which have a similar structure to reward models. Eisenstein (2022) extend and clarify competing notions of ‘feature spuriousness’, using a toy example to show that ‘counterfactual invariance’ and ‘marginally uninformative features’ are distinct notions. Other work examines the generalization of text classifiers when trained on datasets with varying degrees of spurious correlation in features (Kaushik et al., 2020). Joshi et al. (2022) show that spurious correlations can be categorized as ‘necessary’ or ‘sufficient’ for text classifier behavior, and that many ‘necessary but not sufficient’ features interact with other features to affect classifier behavior. RATE extends insights about text classifiers to reward models, which are similarly susceptible to spurious correlations.
Feder et al. (2021) introduced CausaLM, which focuses on training text classifiers to ‘forget’ concepts in order to estimate the treatment effect of an attribute on classification with rule-based rewrites. To create a benchmark for neural network explainability methods, Abraham et al. (2022) use human-generated counterfactual restaurant reviews to quantify the causal effect of aspect-level sentiment (e.g., whether the ambiance was described positively or negatively) on the sentence-level sentiment as predicted by a neural network. Though similar in spirit, RATE is more scalable, as it does not require human-generated counterfactuals, allowing for a diversity of attributes and datasets.
@@
Wang et al. (2024c) survey recent methods for generating counterfactuals. Within their taxonomy, RATE is fully autonomous (using LLM prompting, without using humans to identify the words which need to be changed); furthermore, we offer a stronger evaluation of rewrite quality by rewriting different attributes of the same responses in our synthetic experiments: the fact that RATE is robust to distributional shift validates that rewrites are not changing the secondary off-target concept. Like us, Gat et al. (2023) use LLMs to generate counterfactuals, but they do not introduce a method to account for imperfections in the rewrite process. Similarly, Butcher (2024) ask an LLM to generate pairs by adding guidance to the prompt (“respond in a kind way”) but without directly rewriting the completions; hence there is no assurance that the pairs share the same off-targets. They use datasets with consistent example structure that make it easy to create counterfactual pairs using templates. Wu et al. (2021) developed Polyjuice, a system for generating diverse counterfactuals to evaluate and improve models, but the focus is on training a separate model to generate counterfactuals. Fryer et al. (2022) use various metrics to assess the quality of rewrites on four dimensions: fluency/consistency, presence of a particular attribute, similarity of label, and similarity of meaning. Our work extends assessments of rewrite quality (through rewrites of rewrites) to correct for bias in the evaluation of reward models, allowing us to account for the quality of rewrites on all dimensions simultaneously.
8 Conclusion

### v2

@@ 4 Cancellation of Multiple Rewrite Errors
Intuitively, this assumption is saying that at the level of the reward model there aren’t any interactions between the component corresponding to potential rewrite-errors $\xi$ and either $W$ or $Z$.
The following result establishes validity of RATE under these assumptions:
{restatable}
[Unbiasedness and Consistency of RATE]theoremmainthm Assume $R(\cdot,\cdot)$ is bounded. Take assumptions 1 and 2 above. Suppose we have a set of prompt-completion pairs $\{x^{i},y^{i}\}$ sampled i.i.d. from some population with $P(W=1)\in(0,1)$. Then Algorithm 1 yields unbiased and $\sqrt{n}$-consistent estimators of the ATT, ATU, and ATE.
See Appendix B for the proof.
Original Rewrite Rewrite of Rewrite
It wsa great to see smoe of my favorite stasr of 30 years ago including oJhn Ritter, Ben Gazarra nad Audrye Hepburn. Tehy looked quite wonderful. Great to see some of my favorite stars of 30 years ago, including John Ritter, Ben Gazarra, and Audrey Hepburn. They looked quite wonderful. It was great to see some of my favorite stars of 30 years ago, including John Ritter, Ben Gazarra, and Audrey Hepburn. They looked quite wonderful.
W = 1, Reward: -7.6 W = 0, Reward: -5.6 W = 1, Reward: -5.6
My girlfriend once brought around The Zombie Chronicles for us to watch as a joke. Little did we realize the joke was on her for paying £1 for it. While watching this film I started to come up with things I would rather be doing than watching The Zombie Chronicles. An ex-girlfriend once brought around The Zombie Chronicles for us to watch as a joke. Little did we realize the joke was on her for paying £1 for it. While watching this film I started to come up with things I would rather be doing than watching The Zombie Chronicles. My ex-girlfriend once brought around The Zombie Chronicles for us to watch as a joke. Little did we realize the joke was on her for paying £1 for it. While watching this film I started to come up with things I would rather be doing than watching The Zombie Chronicles.
W = 0, Reward: -5.2 W = 1, Reward: -5.2 W = 0, Reward: -5.1
Table 2: We introduce a correlation between whether the example “starts with a vowel” ($W=1$) and the percentage of words with typos to test double vs single rewrites. In this example, we have added typos to 30% of words, but only to IMDB reviews that start with a vowel. In row one, we see that the original example that starts with a vowel and has typos is penalized by the reward model, while the rewrite of the rewrite has typos fixed and its reward score is the same as the rewrite. Examples that do not start with a vowel ($W=0$) do not have typos introduced and the reward score is similar for the original, the rewrite, and the rewrite of the rewrite. This correlation introduces a spurious positive change in reward when rewriting from “starts with a vowel” to “doesn’t start with a vowel,” yielding a negatively biased estimator for the causal effect of “starts with a vowel” on reward score. Results as the percent of typos is varied are reported in Figure 3.
5 Experiments
There are two main questions to address empirically:
  1. 1.
    Does RATE correctly estimate the causal effect of attributes on reward models?
@@ 5 Experiments
    Is the distinction between RATE and the naive estimator actually substantive?
Answering the first question requires knowing ground truth causal effects. To this end, we design semi-synthetic experiments with known ground truth. Using these, we find that RATE is effective at estimating the true effects, while the naive and single-rewrite estimators fail. The second question is whether the correctness of the RATE estimator over the naive estimator actually matters in practice. We find that across a variety of reward models, attributes, and datasets, the RATE estimates differ substantially from the naive baseline (see Figure 5). Accordingly, the distinction between RATE and the naive estimator is indeed important. As a particular example, the RATE estimator shows that the phenomenon of “length bias” in rewards may actually be mainly an artifact of bias in naive evaluations.
Setup
@@
For all experiments, we use OpenAI BatchAPI to generate rewrites of text, instructing the LLM to modify the target attribute without changing any other aspects of the response (see Section E.1). We use the gpt-4o-2024-08-06 model, incurring a cost of \$1.25 per 1M input tokens and \$5.00 per 1M output tokens. For instance, generating rewrites and rewrites-of-rewrites for 25K IMDB samples cost roughly \$60. See Appendix E for additional implementation details and rewrite samples.
5.1 Semi-synthetic Experiments
Correlation between Typos and “Starts with a Vowel”
To test the efficacy of RATE against a known ground truth, we design a synthetic experiment where we introduce typos into IMDB review examples (Maas et al., 2011) that correlate with whether or not the review starts with a vowel (see Table 2). We score the examples using FsfairX-LLaMA3-RM-v0.1 (Dong et al., 2023) with the prompt: “Write a movie review: ”. We expect the true ATE of “starts with a vowel” on the reward model to be near zero. Additionally, the true effect should not interact with whether there are typos in the review.
Figure 3: In a situation where the ground truth is known, RATE (orange) accurately estimates the ground truth while the naive (blue) and single-rewrite (green) estimators do not. We calculate the treatment effect of “Starts with a vowel” on FsfairX-LLaMA3-RM-v0.1, when typos have been added with varying frequency to IMDB reviews which start with vowels (see Table 2). Intuitively, we expect the RM to respond negatively to the presence of typos, but not to respond at all to whether the movie review starts with a vowel. Hence the treatment effect for ”starts with vowel” should remain zero, even as the spurious correlation increases. 95% confidence intervals are shown.
The naive estimator (Section 2) uses only the original responses; the single-rewrite estimator (Section 3) uses (original, rewrite) pairs, while the RATE estimator (Algorithm 1) uses (rewrite, rewritten rewrite) pairs.
As seen in Table 2, GPT-4o corrects typos even when not asked to, which we predict introduces bias for the single-rewrite estimator. By artificially correlating typos with “starts with a vowel” in the dataset, the single-rewrite procedure ought to demonstrate a positive bias to the change in reward score when examples are rewritten to start with a consonant (since typos harm the reward score). This results in a negatively biased estimate of the treatment effect of “starts with a vowel”—the estimation procedure views “starts with a vowel” as harming the reward score of an example.
In Figure 3, we see that this bias does in fact occur, and increases for both the single rewrite method and the naive estimator as we introduce a higher percentage of typos in examples that start with a vowel. On the other hand, the double rewrite method reports a near-zero treatment effect, even as we increase the percentage of typos in the “starts with a vowel” reviews. This demonstrates the necessity of RATE’s double-rewrite correction.
Treatment Effect of Length on a Sentiment Classifier
As a second test of RATE, we use a DistilBERT sentiment classifier (Socher et al., 2013; Sanh et al., 2020) as a “reward model” since it has the same structure of taking in text and returning a scalar. The benefit of using a sentiment classifier is that it should only be sensitive to the sentiment of the text, not other attributes (e.g. length). Intuitively, the average treatment effect of length corresponds to asking “how much do longer responses impact the likelihood that the DistilBERT model classifies the review as having positive sentiment?” We expect this to be close to zero (assuming no strong correlation between length and sentiment in the classifier’s training data); regardless, the true ATE should be invariant to distributional shift as we increase the correlation between length and positive sentiment.
We induce this correlation by partitioning the IMDB dataset (Maas et al., 2011) into four categories: long positive, short positive, long negative, and short negative reviews. We then downsample each category, keeping the total number of samples constant ($n=9374$) while increasing the correlation between length and positive sentiment (see Table 3 in Section C.2). We then compute the naive, single-rewrite, and RATE estimators on each of the resulting datasets.
Figure 4: Treating a sentiment classifier as a “reward model” gives us an approximate ground truth for the effect of length on sentiment classification. We see that the naive estimator (blue) is again highly sensitive to distributional shift, while both the single-rewrite (green) and RATE (orange) estimators correctly report near zero effects. RATE remains invariant to distributional shift, while the single-rewrite estimator reports an increasingly negative effect as the correlation between length and positive sentiment increases. 95% confidence intervals are shown.
The naive estimator is highly responsive to spurious correlations, while both the single-rewrite estimator and RATE report near-zero treatment effects even in the regime of strong correlation (Figure 4). Noteably, RATE remains invariant to distributional shift, while the single-rewrite estimator reports an increasingly negative effect as the correlation between length and positive sentiment increases.
If we instead replace the sentiment classifer with a general reward model, we can perform a similar test for invariance to distributional shift, just without knowing the ground truth effect. In Section C.1, we reproduce similar results while varying the correlation between complexity and helpfulness in the HelpSteer dataset (Wang et al., 2023). We find that the naive estimator is highly sensitive to distributional shift, while the RATE estimator remains invariant.
Figure 5: An attribute’s reported effect on a reward model differs substantially between the naive (blue) estimator compared to the RATE (orange) estimator. Across reward models, the naive estimator yields much larger effect estimates for length, complexity, and helpfulness; and smaller effect estimates for sentiment. Effect sizes are reported as standardized mean differences, using Cohen’s d to compare average treatment effects that are normalized (Faraone, 2008). Bars represent a 95% confidence interval.
5.2 Real World Reward Models
We select several of the top-performing reward models from RewardBench (Lambert et al., 2024) and evaluate them using both RATE and the naive method across a variety of attributes and datasets: IMDB (Maas et al., 2011), ELI5 (Fan et al., 2019), HelpSteer (Wang et al., 2023).
Figure 5 shows the estimated response of each reward model to each attribute. Of particular interest are the evaluations of FsfairX-LLaMA3-RM-v0.1 (Dong et al., 2023) and NCSOFT (Park et al., 2024) with respect to length. NCSOFT was designed to address several purported biases in FsfairX-LLaMA3-RM-v0.1, including length; however, the RATE estimator suggests that NCSOFT’s improvement here is less than appears at first glance, and may have inadvertently penalized other attributes like complexity.
@@ 6 Related Work
Explainability of Reward Models
Our work is particularly motivated by the challenge of understanding reward model behavior. Jiang et al. (2024) perturb text to find changes that will flip a reward model’s prediction of the “preferred” text example when doing pairwise comparisons—this example-wise approach is complementary to RATE. Lambert et al. (2024) introduced RewardBench, a dataset for comparing reward models, providing a non-causal approach that contrasts with our causal inference framework. Casper et al. (2023); Pan et al. (2022); Tien et al. (2023) highlight issues such as misgeneralization, reward hacking, and spurious associations in reward models, providing motivation for the evaluation developed here.
There is also a directly relevant earlier line of work on spuriousness and interpretability of text classifiers (not necessarily viewed as reward models). Joshi et al. (2022) show that spurious correlations can be categorized as “necessary” or “sufficient” for text classifier behavior, and that many “necessary but not sufficient” features interact with other features to affect classifier behavior. Feder et al. (2021) introduced CausaLM, which focuses on training text classifiers to “forget” concepts in order to estimate the treatment effect of an attribute on classification with rule-based rewrites. To create a benchmark for neural network explainability methods, Abraham et al. (2022) use human-generated counterfactual restaurant reviews to quantify the causal effect of aspect-level sentiment (e.g., whether the ambiance was described positively or negatively) on the sentence-level sentiment as predicted by a neural network. RATE may be seen as a generalization of these insights, using the double-rewrite to allow scaling the creation of counterfactual pairs.
@@
Wang et al. (2024) survey recent methods for generating counterfactuals. Like us, Gat et al. (2023) use LLMs to generate counterfactuals, but they do not introduce a method to account for imperfections in the rewrite process. Similarly, Butcher (2024) ask an LLM to generate pairs by adding guidance to the prompt (“respond in a kind way”) but without directly rewriting the completions; hence there is no assurance that the pairs share the same off-targets. Wu et al. (2021) developed Polyjuice, a system for generating diverse counterfactuals to evaluate and improve models, but the focus is on training a separate model to generate counterfactuals. Fryer et al. (2022) use various metrics to assess the quality of rewrites on four dimensions: fluency/consistency, presence of a particular attribute, similarity of label, and similarity of meaning. Our work extends assessments of rewrite quality (through rewrites of rewrites) to correct for bias in the evaluation of reward models, allowing us to account for the quality of rewrites on all dimensions simultaneously. Bhattacharjee et al. (2024) also make the observation that frontier models are capable of zero-shot generation of counterfactual text examples capable of flipping the label of a text classifier. They use these capabilities to both interpret and create robustness tests for text classifiers, while we rewrite examples on a specific “attribute” and look for a treatment effect of that attribute on a reward model score.
7 Conclusion and Discussion